OpenStudy (anonymous):

Rewrite as an exponential equation. In(x+y)=5

OpenStudy (anonymous):

so ln(x+y)=5 means:\[\log_{e}(x+y)=5 \] so; \[e^5=(x+y)\] is the answer.

OpenStudy (anonymous):

or if you want it terms of y:\[y=e^5-x\]